# Cross-Species Data Integration for Enhanced Layer Segmentation in Kidney Pathology

### v1

@@ 1 INTRODUCTION
In recent years, data-driven deep-learning methods[7, 8] have introduced new tools for clinical computer-aided diagnosis, significantly improving diagnostic efficiency and accuracy. The popularity of convolutional neural networks (CNNs) and Transformer architectures has contributed to a variety of clinical practices, including the segmentation of retinal fundus images [9] and the diagnosis of breast cancer[10]. In the field of kidney diseases, deep-learning models have been employed to identify and segment glomerular structures in human kidney biopsies. However, training high-quality neural network models demands a substantial amount of high-quality annotated data. In the medical field, data collection from humans is particularly challenging due to medical ethics reviews and patient privacy concerns.

### v2

@@ 1 INTRODUCTION
In recent years, data-driven deep-learning methods[7, 8, 9] have introduced new tools for clinical computer-aided diagnosis, significantly improving diagnostic efficiency and accuracy. The popularity of convolutional neural networks (CNNs) and Transformer architectures has contributed to a variety of clinical practices[10, 11], including the segmentation of retinal fundus images [12] and the diagnosis of breast cancer[13]. In the field of kidney diseases, deep-learning models have been employed to identify and segment glomerular structures in human kidney biopsies. However, training high-quality neural network models demands a substantial amount of high-quality annotated data. In the medical field, data collection from humans is particularly challenging due to medical ethics reviews and patient privacy concerns.